OpenStudy (anonymous):

Given the parent functions f(x) = log2 (3x − 9) and g(x) = log2 (x − 3), what is f(x) − g(x)? f(x) − g(x) = log2 (2x − 6) f(x) − g(x) = log2 (2x − 12) f(x) − g(x) = log2 one third f(x) − g(x) = log2 3

OpenStudy (tkhunny):

Have you considered performing the subtraction and using some logarithm rules?

OpenStudy (campbell_st):

well f(x) - g(x) = log2(3x -9) - log2(x -3) the log law for division is the log of the numerator - log of the denominator so you can rewrite it as \[f(x) - g(x) = \log_{2}(\frac{3x -9}{x -3})\] now factor the numerator \[f(x) - g(x) = \log_{2}\frac{3(x -3)}{(x -3)}\] you should be able to simplify it from here.

OpenStudy (anonymous):

didnt work

OpenStudy (tkhunny):

Non Responsive. Please show your work! There is no substitute for knowing what you are doing. Don't just type in an answer. Make sure it is RIGHT. THEN type it in.

OpenStudy (campbell_st):

ok... so latex isn't working well the log law for division is subtract log(a/b) = log(a) - log(b) applying this law to your question its, and the logs are base 2.... \[\log[(3x -9)/(x -3)]\] so all you need to do is factor the numerator and then simplify... you'll get an answer choice from your list hope it helps

OpenStudy (campbell_st):

oh, I'm very reluctant to give answers... I'm a great believer in helping to improve understanding.. so its a little advice fro me... and a little understanding and application from you